# Boundary of the Moduli Space of Stable Cubic Fivefolds

### v6

@@ 4.1 Case $k=1$
Since we work projectively, $\tau$ stabilizes $\varphi_{1}^{\mathrm{nf}}]\in\mathbb{P}(W^{H})$ if and only if $\tau\cdot\varphi_{1}^{\mathrm{nf}}=c\,\varphi_{1}^{\mathrm{nf}}$ for some $c\in\mathbb{C}^{\times}$, equivalently
@@
$\mathrm{Stab}_{T}\!\bigl([\varphi_{1}^{\mathrm{nf}}]\bigr)=H\cdot\{\delta_{\xi}:\ \xi^{7}=1\}\ \cong\ H\times\mu_{7},$
@@ 4.3 Case $k=3$
Write the $x_{1}$-part as a binary quadratic $Q=a_{1}x_{3}^{2}+a_{2}x_{3}x_{4}+a_{3}x_{4}^{2}\in\mathrm{Sym}^{2}\langle x_{3},x_{4}\rangle$. As $\Delta(Q)$ is $\mathrm{SL}(2)$-invariant while rescaling $x_{1}$ scales $Q$ (and hence $\Delta$) homogenously, for a generic (nondegenerate) $Q$, there exists $A\in\mathrm{SL}(2)$ and a rescaling of $x_{1}$ such that $Q\sim x_{3}^{2}+x_{3}x_{4}+x_{4}^{2}$. Thus, the block $\mathrm{(I)}$ is fixed to $x_{1}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{1}x_{4}^{2}$.
@@ 4.9 Case $k=9$
$\varphi^{\mathrm{nf}}_{9}(\tau,\rho)=x_{2}^{2}x_{3}+\tau\,x_{2}x_{3}^{2}+x_{0}x_{4}^{2}+\rho\,x_{0}x_{5}^{2}+x_{1}x_{4}^{2}+x_{1}x_{5}^{2}+x_{0}x_{2}x_{6}+x_{1}x_{3}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$
@@ 4.10 Case $k=10$
$\varphi^{\mathrm{nf}}_{10}(\tau,\rho)=x_{2}^{3}+\tau\,x_{2}^{2}x_{3}+\rho\,x_{2}x_{3}^{2}+x_{3}^{3}+x_{1}x_{2}x_{5}+x_{1}x_{3}x_{4}+x_{0}x_{4}x_{5}+x_{0}x_{6}^{2},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$
@@ 4.13 Case $k=13$
$\varphi^{\mathrm{nf}}_{13}(\rho,\sigma)=x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{5}^{2}+x_{1}x_{3}x_{4}+\rho\,x_{2}x_{3}x_{5}+x_{1}^{2}x_{6}+\sigma\,x_{2}^{2}x_{6}+x_{0}x_{3}x_{6},\qquad(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}.$
@@ 4.14 Case $k=14$
$\varphi^{\mathrm{nf}}_{14}(\tau)=x_{2}^{3}+x_{0}\bigl(x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)+x_{1}\bigl(x_{3}^{2}+\tau x_{4}^{2}-x_{5}^{2}+2x_{6}^{2}\bigr),\qquad\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}.$
@@ 4.15 Case $k=15$
$\varphi^{\mathrm{nf}}_{15}(\tau,\rho)\;=\;x_{3}^{2}x_{4}\;+\;\tau\,x_{3}x_{4}^{2}\;+\;x_{0}x_{3}x_{5}\;+\;x_{0}x_{4}x_{6}\;+\;x_{1}^{2}x_{5}\;+\;\rho\,x_{2}^{2}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$ (20)
@@ 4.18 Case $k=18$
$\varphi^{\mathrm{nf}}_{18}=x_{2}^{3}+\sigma_{1}x_{3}^{3}+\sigma_{2}x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+x_{0}x_{2}x_{5}+x_{1}x_{3}x_{5}+\alpha\,x_{0}x_{4}x_{5}+x_{0}x_{2}x_{6}+\beta\,x_{1}x_{3}x_{6}+\gamma\,x_{1}x_{4}x_{6},$
@@ 4.19 Case $k=19$
$\phi_{19}^{\mathrm{nf}}(\rho,\sigma)=x_{3}^{3}+(x_{0}x_{4}^{2}+x_{1}x_{4}x_{5}+x_{2}x_{5}^{2})+x_{6}(x_{0}^{2}+x_{2}^{2}+\rho\,x_{1}^{2}+\sigma\,x_{0}x_{2}),$
@@ 4.20 Case $k=20$
$\varphi^{\mathrm{nf}}_{20}(\tau)=x_{4}^{3}+x_{5}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})+x_{6}(x_{0}^{2}+\tau x_{1}^{2}-x_{2}^{2}+2x_{3}^{2}),\qquad\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}.$
@@ 4.21 Case $k=21$
where $\sum v_{i}=0$, and with the $\mathrm{SL}$-constraint $S:=2\alpha+4\beta+\gamma=0$. The weight of any monomial in $F$ is $3\beta$, and the weight of any monomial in $x_{6}Q$ is $\gamma+2\alpha$. Using $S=0$, we have $\gamma+2\alpha=-4\beta$. Thus we satisfy the positive linear identity
$3\,w(x_{6}Q)\;+\;4\,w(F)\;=\;3(-4\beta)+4(3\beta)\;=\;0.$ (27)
If $\lambda\in\Lambda_{\phi_{21}}$, weights are non-negative, so they must vanish: $\beta=0$ (implies $v_{i}=0$) and $\gamma+2\alpha=0$. The condition $w(x_{6}Q)=0$ with $u=0$ (from generic stability of $Q$) implies $\lambda$ reduces to the form of $\lambda_{21}$ (up to scaling). Thus $\Lambda_{\phi_{21}}$ is symmetric, and by the Casimiro–Florentino criterion, $\phi_{21}$ is polystable.
@@
14 C-F $x_{2}^{3}+x_{0}\bigl(x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)+x_{1}\bigl(x_{3}^{2}+\tau x_{4}^{2}-x_{5}^{2}+2x_{6}^{2}\bigr)$ $\tau\in\mathbb{C}^{\times}$, $\tau\neq\pm 1,2$ 11
@@ 5 Singular loci of 21 polystable cubic fivefolds
In this section, we determine the singular loci of the closed–orbit representatives $\phi^{\mathrm{nf}}_{k}$ constructed in Section 4. For each $k=1,\dots,21$, we set
@@
i.e. the scheme cut out by the saturated Jacobian ideal. We give a set–theoretic description of $\mathrm{Sing}(X_{k})$ in each case, and when isolated points occur, we also record the corank and the local invariants (Milnor and Tjurina numbers, which agree for our quasi–homogeneous normal forms). Table 3 presents a compact summary—listing the type and degree of the top–dimensional part and indicating the presence (or absence) of isolated points. Detailed case–by–case statements are recorded as Propositions 5.85.28.
@@
We define the isolated singularities that appear in the investigation of the singularities of $\varphi_{k}^{\mathrm{nf}}$. Since we could not locate an explicit normal form in the standard tables (e.g. [AGV12]), so we fix notation here.
@@
We write $\mathrm{QH}(9,6,5;24)_{19}$ for the isolated hypersurface singularity analytically equivalent to $X^{2}Y+Y^{4}+XZ^{3}$. It is quasi-homogeneous of total degree $24$ with respect to weights $(w_{X},w_{Y},w_{Z})=(9,6,5)$. Its corank, Milnor number and Tjurina numbers are 33,1919 and $19$ respectively.
@@
We write $\mathrm{QH}(2,3,5;12)_{21}(t)$ for the isolated hypersurface singularity analytically equivalent to $tX^{6}+X^{3}Y^{2}+X^{2}YZ+XZ^{2}+Y^{4}$, where $t\in\mathbb{C}^{\times}$ is a parameter. It is quasi-homogeneous of total degree $12$ with respect to weights $(w_{X},w_{Y},w_{Z})=(2,3,5)$. Its corank, Milnor number and Tjurina numbers are 33,2121 and $21$ respectively.
@@
We write $\mathrm{QH}(3,3,4;12)_{18}(t)$ for the isolated hypersurface singularity analytically equivalent to $Z^{3}+X^{4}+tX^{2}Y^{2}+Y^{4}$. It is quasi-homogeous of total degree $12$ with respect to weights $(w_{X},w_{Y},w_{Z})=(3,3,4)$. Its corank, Milnor number and Tjurina number are $3$, $18$ and $18$ respectively.
@@ 5.2 Case $k=2$
Let $X_{2}=V(\varphi^{\mathrm{nf}}_{2})\subset\mathbb{P}^{6}$. Then
@@ 5.3 Case $k=3$
Let $X_{3}=V(\varphi^{\mathrm{nf}}_{3})\subset\mathbb{P}^{6}$. Then
@@ 5.4 Case $k=4$
Let $X_{4}=V(\varphi^{\mathrm{nf}}_{4})\subset\mathbb{P}^{6}$. Then
@@ 5.5 Case $k=5$
Let $X_{5}=V(\varphi^{\mathrm{nf}}_{5})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{5})=P\cup Q$
$P=(1:0:0:0:0:0:0)$, $Q=(0:0:0:0:0:0:1)$ are isolated singular points.
@@ 5.6 Case $k=6$
Let $X_{6}=V(\varphi^{\mathrm{nf}}_{6})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{6})=\{P\}\cup V(x_{3},x_{4},x_{5},x_{6},\;x_{1}^{2}+x_{0}x_{2}),\quad P=(0:0:0:0:0:0:1).$
The isolated singular point $P$ is of type $\mathsf{E}_{19}$.
@@ 5.7 Case $k=7$
Let $X_{7}=V(\varphi^{\mathrm{nf}}_{7})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{7})=C\cup L_{56},\quad L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4}),$
$C$ is a $\mathrm{CI}(2,2)$ of degree $4$; $C\cap L_{56}=\varnothing$.
@@ 5.13 Case $k=13$
Let $X_{13}:=V(\phi^{\mathrm{nf}}_{13})\subset\mathbb{P}^{6}$. Then the singular locus has two connected components:
@@
where $C_{1}=V(x_{1}^{2}+3x_{2}^{2},x_{3},x_{4},x_{5},x_{6})$ and $C_{2}=V(x_{0},x_{1},x_{2},x_{3},x_{4}^{2}+x_{5}^{2})$. Each $C_{i}$ consists of two lines which intersect at one point. We have $C_{1}\cap C_{2}=\varnothing$.

### v7

@@ 4.1 Case $k=1$
Since we work projectively, $\tau$ stabilizes $\phi^{\mathrm{nf}}_{1}]\in\mathbb{P}(W^{H})$ if and only if $\tau\cdot\phi^{\mathrm{nf}}_{1}=c\,\phi^{\mathrm{nf}}_{1}$ for some $c\in\mathbb{C}^{\times}$, equivalently
@@
$\mathrm{Stab}_{T}\!\bigl([\phi^{\mathrm{nf}}_{1}]\bigr)=H\cdot\{\delta_{\xi}:\ \xi^{7}=1\}\ \cong\ H\times\mu_{7},$
@@ 4.3 Case $k=3$
Write the $x_{1}$-part as a binary quadratic $Q=a_{1}x_{3}^{2}+a_{2}x_{3}x_{4}+a_{3}x_{4}^{2}\in\mathrm{Sym}^{2}\langle x_{3},x_{4}\rangle$. As $\Delta(Q)$ is $\mathrm{SL}(2)$-invariant while rescaling $x_{1}$ scales $Q$ (and hence $\Delta$) homogeneously, for a generic (nondegenerate) $Q$, there exists $A\in\mathrm{SL}(2)$ and a rescaling of $x_{1}$ such that $Q\sim x_{3}^{2}+x_{3}x_{4}+x_{4}^{2}$. Thus, the block $\mathrm{(I)}$ is fixed to $x_{1}x_{3}^{2}+x_{1}x_{3}x_{4}+x_{1}x_{4}^{2}$.
@@ 4.9 Case $k=9$
$\phi^{\mathrm{nf}}_{9}(\tau,\rho)=x_{2}^{2}x_{3}+\tau\,x_{2}x_{3}^{2}+x_{0}x_{4}^{2}+\rho\,x_{0}x_{5}^{2}+x_{1}x_{4}^{2}+x_{1}x_{5}^{2}+x_{0}x_{2}x_{6}+x_{1}x_{3}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$
@@ 4.10 Case $k=10$
$\phi^{\mathrm{nf}}_{10}(\tau,\rho)=x_{2}^{3}+\tau\,x_{2}^{2}x_{3}+\rho\,x_{2}x_{3}^{2}+x_{3}^{3}+x_{1}x_{2}x_{5}+x_{1}x_{3}x_{4}+x_{0}x_{4}x_{5}+x_{0}x_{6}^{2},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}$
@@ 4.13 Case $k=13$
$\phi^{\mathrm{nf}}_{13}(\rho,\sigma)=x_{3}^{3}+x_{0}x_{4}^{2}+x_{0}x_{5}^{2}+x_{1}x_{3}x_{4}+\rho\,x_{2}x_{3}x_{5}+x_{1}^{2}x_{6}+\sigma\,x_{2}^{2}x_{6}+x_{0}x_{3}x_{6},\qquad(\rho,\sigma)\in(\mathbb{C}^{\times})^{2}.$
@@ 4.14 Case $k=14$
$\phi^{\mathrm{nf}}_{14}(\tau)=x_{2}^{3}+x_{0}\bigl(x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)+x_{1}\bigl(x_{3}^{2}+\tau x_{4}^{2}-x_{5}^{2}+2x_{6}^{2}\bigr),\qquad\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}.$
@@ 4.15 Case $k=15$
$\phi^{\mathrm{nf}}_{15}(\tau,\rho)\;=\;x_{3}^{2}x_{4}\;+\;\tau\,x_{3}x_{4}^{2}\;+\;x_{0}x_{3}x_{5}\;+\;x_{0}x_{4}x_{6}\;+\;x_{1}^{2}x_{5}\;+\;\rho\,x_{2}^{2}x_{6},\qquad(\tau,\rho)\in(\mathbb{C}^{\times})^{2}.$ (20)
@@ 4.18 Case $k=18$
$\phi^{\mathrm{nf}}_{18}=x_{2}^{3}+\sigma_{1}x_{3}^{3}+\sigma_{2}x_{4}^{3}+\rho\,x_{2}x_{3}x_{4}+x_{0}x_{2}x_{5}+x_{1}x_{3}x_{5}+\alpha\,x_{0}x_{4}x_{5}+x_{0}x_{2}x_{6}+\beta\,x_{1}x_{3}x_{6}+\gamma\,x_{1}x_{4}x_{6},$
@@ 4.19 Case $k=19$
$\phi^{\mathrm{nf}}_{19}(\rho,\sigma)=x_{3}^{3}+(x_{0}x_{4}^{2}+x_{1}x_{4}x_{5}+x_{2}x_{5}^{2})+x_{6}(x_{0}^{2}+x_{2}^{2}+\rho\,x_{1}^{2}+\sigma\,x_{0}x_{2}),$
@@ 4.20 Case $k=20$
$\phi^{\mathrm{nf}}_{20}(\tau)=x_{4}^{3}+x_{5}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})+x_{6}(x_{0}^{2}+\tau x_{1}^{2}-x_{2}^{2}+2x_{3}^{2}),\qquad\tau\in\mathbb{C}^{\times}\setminus\{1,-1,2\}.$
@@ 4.21 Case $k=21$
where $\sum v_{i}=0$, and with the $\mathrm{SL}$-constraint $S:=2\alpha+4\beta+\gamma=0$.
Summing the $\lambda$-weights of the $20$ cubic monomials in $F$ gives
$w(F)=15\bigl((\beta+v_{1})+(\beta+v_{2})+(\beta+v_{3})+(\beta+v_{4})\bigr)=60\beta,$
while the weights of the three quadratic monomials in $x_{6}Q$ sum to
$w(x_{6}Q)=(\gamma+2(\alpha+u))+(\gamma+2\alpha)+(\gamma+2(\alpha-u))=3\gamma+6\alpha.$
Using $S=0$, we obtain the positive linear identity
$w(F)+5\,w(x_{6}Q)\;=\;60\beta+5(3\gamma+6\alpha)\;=\;15(2\alpha+4\beta+\gamma)\;=\;15\,S.$ (27)
If $\lambda\in\Lambda_{\phi_{21}}$, then every monomial weight occurring in $F$ and $x_{6}Q$ is $\geq 0$ and $S=0$; hence (27) forces
$w(F)=0,\qquad w(x_{6}Q)=0,$
so all those monomial weights vanish individually. From the cubic monomials $x_{2}^{3},\dots,x_{5}^{3}$ we get $\beta+v_{i}=0$ for all $i$, hence $\beta=0$ and $v_{i}=0$. From the quadratic monomials $x_{6}x_{0}^{2}$ and $x_{6}x_{1}^{2}$ we get
$\gamma+2\alpha+2u=0,\qquad\gamma+2\alpha-2u=0,$
hence $u=0$ and $\gamma+2\alpha=0$. Therefore $\lambda$ is of the form
$\lambda(t)=\operatorname{diag}\!\bigl(t^{k},\,t^{k},\,1,\,1,\,1,\,1,\,t^{-2k}\bigr)=\lambda_{21}(t^{k})\qquad(k\in\mathbb{Z}),$
and consequently
$\Lambda_{\phi_{21}}=\{\lambda_{21}(t^{k})\mid k\in\mathbb{Z}\}.$
By Lemma 4.5, $\Lambda_{\phi_{21}}$ is symmetric; hence, by the Casimiro–Florentino criterion, $\phi_{21}$ is polystable.
@@
The dimension of the component is computed as follows: $\dim W^{H}=20(\text{cubic})+3(\text{quadratic})=23$. Since $H$ acts trivially on $W^{H}$, the effective group is $C_{G}(H)/H$, of dimension 201=1920-1=19. Since the generic stabilizer has dimension $1$ (due to $\mathrm{O}(2)$), the orbit dimension is 191=1819-1=18.
@@
14 C-F $x_{2}^{3}+x_{0}\bigl(x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}\bigr)+x_{1}\bigl(x_{3}^{2}+\tau x_{4}^{2}-x_{5}^{2}+2x_{6}^{2}\bigr)$ $\tau\in\mathbb{C}^{\times}\setminus\{-1,1,2\}$ 11
@@ 5 Singular loci of 21 polystable cubic fivefolds
In this section, we determine the singular loci of general closed–orbit representatives $\phi^{\mathrm{nf}}_{k}$ constructed in Section 4. For each $k=1,\dots,21$, we set
@@
i.e. the scheme cut out by the saturated Jacobian ideal. We give a set–theoretic description of $\mathrm{Sing}(X_{k})$ in each case, and when isolated points occur, we also record the corank and the local invariants (Milnor and Tjurina number, which agree for our quasi–homogeneous normal forms). Table 3 presents a compact summary—listing the type and degree of the top–dimensional part and indicating the presence (or absence) of isolated points. Detailed case–by–case statements are recorded as Propositions 5.85.28.
Note that for the boundary families with moduli (parameters), the descriptions of the singular loci in the following propositions and the minimal exponent computations in Section 6 apply to general parameters. At certain special values of the parameters, the singularities may degenerate further, and the minimal exponent may strictly drop.
@@
We define the isolated singularities that appear in the investigation of the singularities of $\phi^{\mathrm{nf}}_{k}$. Since we could not locate an explicit normal form in the standard tables (e.g. [AGV12]), we fix notation here.
@@
We write $\mathrm{QH}(9,6,5;24)_{19}$ for the isolated hypersurface singularity analytically equivalent to $X^{2}Y+Y^{4}+XZ^{3}$. It is quasi-homogeneous of total degree $24$ with respect to weights $(w_{X},w_{Y},w_{Z})=(9,6,5)$. Its corank, Milnor number and Tjurina number are $3$, 1919 and $19$ respectively.
@@
We write $\mathrm{QH}(2,3,5;12)_{21}(t)$ for the isolated hypersurface singularity analytically equivalent to $tX^{6}+X^{3}Y^{2}+X^{2}YZ+XZ^{2}+Y^{4}$, where $t\in\mathbb{C}^{\times}$ is a parameter. It is quasi-homogeneous of total degree $12$ with respect to weights $(w_{X},w_{Y},w_{Z})=(2,3,5)$. Its corank, Milnor number and Tjurina number are $3$, 2121 and $21$ respectively.
@@
We write $\mathrm{QH}(3,3,4;12)_{18}(t)$ for the isolated hypersurface singularity analytically equivalent to $Z^{3}+X^{4}+tX^{2}Y^{2}+Y^{4}$. It is quasi-homogeneous of total degree $12$ with respect to weights $(w_{X},w_{Y},w_{Z})=(3,3,4)$. Its corank, Milnor number and Tjurina number are $3$, $18$ and $18$ respectively.
@@ 5.2 Case $k=2$
Let $X_{2}=V(\phi^{\mathrm{nf}}_{2})\subset\mathbb{P}^{6}$. Then
@@ 5.3 Case $k=3$
Let $X_{3}=V(\phi^{\mathrm{nf}}_{3})\subset\mathbb{P}^{6}$. Then
@@ 5.4 Case $k=4$
Let $X_{4}=V(\phi^{\mathrm{nf}}_{4})\subset\mathbb{P}^{6}$. Then
@@ 5.5 Case $k=5$
Let $X_{5}=V(\phi^{\mathrm{nf}}_{5})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{5})=\{P\}\cup\{Q\},$
where $P=(1:0:0:0:0:0:0)$, $Q=(0:0:0:0:0:0:1)$ are isolated singular points.
@@ 5.6 Case $k=6$
Let $X_{6}=V(\phi^{\mathrm{nf}}_{6})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{6})=C\cup L_{56},\quad L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4}),$
$C$ is a $\mathrm{CI}(2,2)$ of degree $4$; $C\cap L_{56}=\varnothing$.
@@ 5.7 Case $k=7$
Let $X_{7}=V(\phi^{\mathrm{nf}}_{7})\subset\mathbb{P}^{6}$. Then
$\mathrm{Sing}(X_{7})=C\cup\{P\},$
where
$C=V(x_{3},x_{4},x_{5},x_{6},\;x_{1}^{2}+x_{0}x_{2}),\quad P=(0:0:0:0:0:0:1).$
The isolated singular point $P$ is of type $\mathsf{E}_{19}$.
@@ 5.13 Case $k=13$
Let $X_{13}:=V(\phi^{\mathrm{nf}}_{13}(\rho,\sigma))\subset\mathbb{P}^{6}$. Then the singular locus has two connected components:
@@
where $C_{1}=V(x_{1}^{2}+\sigma x_{2}^{2},x_{3},x_{4},x_{5},x_{6})$ and $C_{2}=V(x_{0},x_{1},x_{2},x_{3},x_{4}^{2}+x_{5}^{2})$. Each $C_{i}$ consists of two lines which intersect at one point. We have $C_{1}\cap C_{2}=\varnothing$.